OpenStudy (anonymous):

Anna throws a ball horizontally off her apartment balcony at a speed of 6.0 meters/second. If the ball falls 10 meters from the balcony to the base of the apartment, calculate how long it takes the ball to reach the base. Choices: -0.6 seconds -1.7 seconds -10 seconds -60 seconds I was thinking it's 60 seconds, because I multiplied 6 by 10, but I don't trust myself..

OpenStudy (anonymous):

so, v=x/t => v/x=1/t => x/v=t right?

OpenStudy (anonymous):

o-o I have no idea..

OpenStudy (anonymous):

v=x/t is just a formula you should know

OpenStudy (anonymous):

form this you can know what t is

OpenStudy (anonymous):

which is 1.7

OpenStudy (anonymous):

Oh okay~ I'll plug it in ^-^ Thanks a lot!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

I got it right! Thanks a million. Good thing I didn't trust myself..